OpenStudy (anonymous):

Car 1 and Car 2 have an equal mass of 1,000 kg each. They were involved in a rear-end accident at a stop sign. here is what we know 1. car 1 approched the intersection from the top of a 25 meter hill 2.car 2 was on a flat stretch of road directly in front of car 1 3.at the bottom of the hill, Car 1 was going 20 m/s and Car 2 was going 30m/s before it stopped at the stop sign. 4.there were no sid marks left by car 2. Car 2 stopeed at the stop sign 5.After the acciedent, both cars were moving in the same direction 10m/s 6. we must push car 2 using 800 n of force, 10 meters off the road.

OpenStudy (anonymous):

Thank you so much whoever that is going to help me. I'm doing this pratice worksheet. and I'm in honors biology, so I am totally not a physics person. ThankS!!!1

OpenStudy (youngster):

Wait, what do you want to find?

OpenStudy (anonymous):

how much energy did car 1 lose from the top to the bottom of the hill. Thanks for coming!

OpenStudy (youngster):

okay-so first find the PE for it at the top of the hill. (assuming it was not moving) PE=mgh PE=1,000x9.8x25=245,000 Then find the KE at the bottom of the hill KE=1/2 mv^2 KE=.5x1,000x(20^2)=200,000 Difference=245,000-200,000= 45,000 J # My answer seems awfully large so you should check my work.

OpenStudy (anonymous):

wait so its the difference? I thought is should be a sum?

OpenStudy (anonymous):

i thought it was energy totaly=PE+KE?

OpenStudy (youngster):

It is the PE at the top and the KE at the bottom.

OpenStudy (youngster):

and it is assuming no initial movement

OpenStudy (anonymous):

@Youngster : what abt the energy Car 1 loses during collision?

OpenStudy (anonymous):

I can't seem to fix it in the calculation. there's also the case of point no. 6. Work done in pushing Car2 = Force . displacement = 800 . 10 = 8000N will this be the energy Car1 loses during collision? If that's the case, then the answer would be 2,45,000 - 2,00,000 - 8000 = 37,000 J

OpenStudy (anonymous):

@jokereq7 : do u know the final answer?